OpenStudy (anonymous):

Find the coordinates of three more points that lie on the line passing through the points (0,0) and (3,-4)

OpenStudy (anonymous):

have you worked with equations of lines yet? Like, y = mx + b where m is slope and b is the y-intercept?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I just want to know how to find the points

OpenStudy (anonymous):

Can you find \(y=mx+b\) for the line that goes through these points?

OpenStudy (anonymous):

You first step is to find that line. Then you have to plug in sample x values to get the corresponding y values. Each x, y pair is a coordinate on the same line as those points.

OpenStudy (anonymous):

Can the sample x values be any number?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

who do i find y=mx+b?

OpenStudy (anonymous):

any x value will produce a corresponding y value on that line. It's usually best to choose "easy" x values, but you could pick x=100 if you want

OpenStudy (anonymous):

Can you show me an example, i don't get it

OpenStudy (anonymous):

to get slope, m, you take the difference in the y values divided by the difference in the x values. m = (-4 - 0) / (3 - 0) = - 4 /3

OpenStudy (anonymous):

i know that, then what?

OpenStudy (anonymous):

ok, so the line has slope -4/3 and it passes, thankfully, through the origin, so you already have the y-intercept, b, equal to 0 So the expression for the line is y = -(4/3)x Now pick 3 values of x... multiples of 3 would make it easy... and put in your x to find the corresponding y... those become ordered pair point on the line y = -(4/3)x but don't choose x = 3... you already know (3,-4) is on the line. Try x = 6, x = 9, and x = 12

OpenStudy (anonymous):

so (6, ___) (9, ___) and (12, ___)

OpenStudy (anonymous):

thank you so much! that was a great explanation!

OpenStudy (anonymous):

great :) Glad to help :)

OpenStudy (anonymous):

So then.... y=-4/3(6) y= -24/3 y= -8

OpenStudy (anonymous):

yes, so (6, -8) is a point on the line... you got this now :) Repeat for the other 2 points.